# On von Neumann regularity of ample groupoid algebras

### v2

@@ 2.4. Regular rings
We know that $S=\Gamma_{c}(\mathscr{G})$ is locally finite by Proposition 3.4. We show that no maximal subgroup of $S$ contains an element of order divisible by the characteristic of $K$. It will then follow that $KS$ is regular by Proposition 2.5, and hence its quotient $K\mathscr{G}$ is regular by Proposition 2.1. Let $U\subseteq\mathscr{G}^{(0)}$ be the identity of a maximal subgroup $H$, and note that $H$ consists of those $V\in S$ with $\mathop{\boldsymbol{s}}(V)=U=\mathop{\boldsymbol{r}}(V)$. If $V\in H$ and $x\in U$, then denote by $v_{x}$ the unique element of $V$ with $\mathop{\boldsymbol{r}}(v_{x})=x$. Since $\mathscr{G}$ consists of isotropy, $\mathop{\boldsymbol{s}}(v_{x})=x$ as well. Then we have an injective group homomorphism $\psi\colon H\to\prod_{x\in U}\mathscr{G}_{x}^{x}$ given by $\psi(V)_{x}=v_{x}$. Suppose that $K$ has characteristic $p>0$. Note that if $V\in H$ satisfies $V^{p}=U$, then $\psi(V)^{p}=\psi(U)$ and hence $v_{x}^{p}=x$ for all $x\in U$. Since no isotropy group has an element of order $p$, we deduce that $v_{x}=x$ for all $x\in U$, and hence $V=U$. Thus no element of $H$ has order divisible by $p$. This concludes the proof. ∎
@@
Let $\mathscr{G}$ be a quasi-compact ample groupoid and $K$ a field whose characteristic does not divide the order of any isotropy group element. Suppose that all orbits of $\mathscr{G}$ have the same cardinality. Then $K\mathscr{G}$ is regular.
@@
It follows from Lemma 3.3 that the size of each orbit is finite, say that this common size is $n$. Choose for each $x\in\mathscr{G}^{(0)}$, elements $\gamma_{x,1},\ldots,\gamma_{x,n}$ with $\mathop{\boldsymbol{s}}(\gamma_{x,i})=x$ and $\mathcal{O}_{x}=\{\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x,n})\}$. We may assume without loss of generality that $\gamma_{x,1}=x$ (the identity at $x$). Choose for each $i$ a compact open bisection $U^{x}_{i,1}$ containing $\gamma_{x,i}$. We may assume that $U^{x}_{1,1}\subseteq\mathscr{G}^{(0)}$. Since $\mathscr{G}^{(0)}$ is Hausdorff with a basis of compact open sets and $\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x_{n}})$ are pairwise distinct, we can find pairwise disjoint compact open neighborhoods $V_{1},\ldots,V_{n}$ of $\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x,n})$, respectively, in $\mathscr{G}^{(0)}$. Replacing $U^{x}_{i,1}$ by $V_{i}U^{x}_{i,1}$, we may assume that the sets $U^{x}_{i}=\mathop{\boldsymbol{r}}(U^{x}_{i,1})$ are pairwise disjoint. Also, putting $U=\bigcap_{i=1}^{n}\mathop{\boldsymbol{s}}(U^{x}_{i,1})$ and replacing $U^{x}_{i,1}$ by $U^{x}_{i,1}U$, we may assume that $\mathop{\boldsymbol{s}}(U^{x}_{i,1})=\mathop{\boldsymbol{s}}(U^{x}_{1,1})=U^{x}_{1}$ for all $i$. Put $U^{x}_{i,j}=U^{x}_{i,1}(U^{x}_{j,1})^{-1}$. Note that $\mathop{\boldsymbol{s}}(U^{x}_{i,j})=U^{x}_{j}$ and $\mathop{\boldsymbol{r}}(U^{x}_{i,j})=U^{x}_{i}$ for all $i,j$. Also note that $U^{x}_{i,i}=U^{x}_{i}$ for $i=1,\ldots,n$. Let $U^{x}=\bigcup_{i=1}^{n}U^{x}_{i}$. Then $U^{x}$ is compact open and we claim that it is invariant. For suppose that $y\in U^{x}_{j}$. Then since $U^{x}_{j}=\mathop{\boldsymbol{s}}(U^{x}_{i,j})$ and $U^{x}_{i}=\mathop{\boldsymbol{r}}(U^{x}_{i,j})$ for all $i=1,\ldots,n$, we can find elements of the orbit of $y$ in each of $U^{x}_{1},\ldots,U^{x}_{n}$. Since these sets are pairwise disjoint, this yields $n$ elements in the orbit of $y$ and hence the whole orbit by assumption. Thus each $U^{x}$ is invariant and each $U^{x}_{i}$ contains exactly one element from each orbit of $U^{x}$. Also note that
@@
Another observation is that the isotropy subgroupoid $\mathop{\mathrm{Iso}}(\mathscr{G})$ of $\mathscr{G}$ is clopen, and hence quasi-compact. It is closed since $\mathscr{G}^{(0)}$ is Hausdorff and it is the equalizer of $\mathop{\boldsymbol{s}},\mathop{\boldsymbol{r}}$. It is open because if $\gamma\colon x\to x$ is isotropy and $V$ is a compact open bisection containing $\gamma$, then $U^{x}_{1}VU^{x}_{1}$ is a compact open bisection containing $\gamma$ and consisting of isotropy as $x\in U^{x}_{1}$, and $U^{x}_{1}$ contains exactly only one element from each orbit of the invariant compact open set $U^{x}$, whence $\mathop{\boldsymbol{s}}(\alpha),\mathop{\boldsymbol{r}}(\alpha)\in U^{x}_{1}$ implies $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}(\alpha)$.
Since $\mathscr{G}^{(0)}$ is compact and the $U^{x}$ cover $\mathscr{G}^{(0)}$, we can find finitely may points $x_{1},\ldots,x_{r}$ such that $\mathscr{G}^{(0)}=U^{x_{1}}\cup\cdots\cup U^{x_{r}}$. Without loss of generality, we may assume that $U^{x_{i}}\nsubseteq\bigcup_{j\neq i}U^{x_{j}}$ for all $i$. Put $V_{1}=U^{x_{1}},V_{2}=U^{x_{2}}\setminus V_{1},\ldots,V_{r}=U^{x_{r}}\setminus\bigcup_{i=1}^{r-1}V_{i}$. Then $V_{1},\ldots,V_{r}$ is partition of $\mathscr{G}^{(0)}$ into compact open invariant sets. In particular note that $V_{i}U=UV_{i}$ for any compact open bisection $U$. It follows that if $\mathscr{G}_{i}=\mathscr{G}|_{V_{i}}$, then $K\mathscr{G}\cong\prod_{k=1}^{r}K\mathscr{G}_{k}$ and $K\mathscr{G}_{k}=1_{V_{k}}K\mathscr{G}1_{V_{k}}$.
Note that since $\mathscr{G}_{k}=\mathop{\boldsymbol{r}}^{-1}(V_{k})$ is clopen in $\mathscr{G}$ it is quasi-compact. Also $\mathop{\mathrm{Iso}}(\mathscr{G}_{k})=\mathop{\mathrm{Iso}}(\mathscr{G})\cap\mathscr{G}_{k}$ is clopen and hence quasi-compact. Put $U^{k}_{i,j}=V_{k}U^{x_{k}}_{i,j}=U^{x_{k}}_{i,j}V_{k}$; this is a compact open bisection in $\mathscr{G}_{k}$. Also note that $U^{k}_{1,1}\subseteq U^{x_{k}}_{1}$ does not contain any two points in the same orbit by the previous discussion. Hence $U^{k}_{1,1}\mathscr{G}_{k}U^{k}_{1,1}=\mathscr{G}_{k}|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})\cap\mathop{\boldsymbol{r}}^{-1}(U^{k}_{1,1})$ is clopen (as $U^{k}_{1,1}$ is a clopen subspace of $\mathscr{G}^{(0)}$) and thus a quasi-compact group bundle. Note that $V_{k}=V_{k}U^{x_{k}}=\bigcup_{i=1}^{n}U^{k}_{i,i}$ and hence it follows from (3.2) that the $1_{U^{k}_{i,j}}$ form a set of matrix units for $K\mathscr{G}_{k}$. Thus $K\mathscr{G}_{k}\cong M_{n}(1_{U^{k}_{1,1}}K\mathscr{G}_{k}1_{U^{k}_{1,1}})=M_{n}(K\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}})$. We deduce that $K\mathscr{G}$ is regular from Propositions 2.1 and 3.9. ∎
@@
Clearly $O_{k}$ is invariant; we claim that it is open. Suppose that $x\in O_{k}$. Let $x=y_{1},y_{2}\ldots,y_{k}$ be $k$ distinct elements in the orbit of $x$. As in the proof of Lemma 3.10, we can find open bisections $U_{1},\ldots,U_{k}$ with $x\in\mathop{\boldsymbol{s}}(U_{i})$, $y_{i}\in\mathop{\boldsymbol{r}}(U_{i})$, for $i=1,\ldots,k$, and $\mathop{\boldsymbol{r}}(U_{1}),\ldots,\mathop{\boldsymbol{r}}(U_{k})$ pairwise disjoint (using that $\mathscr{G}^{(0)}$ is Hausdorff). Then $U=\bigcap_{i=1}^{k}\mathop{\boldsymbol{s}}(U_{i})$ is a neighborhood of $x$ and if $y\in U$, then each of $\mathop{\boldsymbol{r}}(U_{1}),\ldots,\mathop{\boldsymbol{r}}(U_{k})$ contains an element of the orbit of $y$. Since these sets are pairwise disjoint, we conclude that $y\in O_{k}$. ∎
We can now prove our second main theorem.
Theorem 3.12.
Suppose that $K$ is a field and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that the order of no isotropy element is divisible by the characteristic of $K$. Then $K\mathscr{G}$ is regular.
@@
Let $d_{1}<d_{2}<\cdots<d_{m}$ be the distinct orbit sizes. We prove the result by induction on $m$. If $m=1$, then all orbits of $\mathscr{G}$ have the same size and the result follows from Lemma 3.10. Let $U=\{x\in\mathscr{G}^{(0)}\mid|\mathcal{O}_{x}|=d_{m}\}$. Then $U$ is open and invariant by Lemma 3.11 once we note that $U=O_{d_{m}}$ in the notation of that lemma. Let $X=\mathscr{G}^{(0)}\setminus U$. Then $I=K\mathscr{G}|_{U}$ is an ideal of $K\mathscr{G}$ and $K\mathscr{G}/I\cong K\mathscr{G}|_{X}$. Note that $\mathscr{G}|_{U}=\mathop{\boldsymbol{r}}^{-1}(U)$ is open and $\mathscr{G}|_{X}=\mathop{\boldsymbol{r}}^{-1}(X)$ is closed and hence quasi-compact. By construction, the orbits sizes in $\mathscr{G}|_{X}$ are $d_{1}<d_{2}<\cdots<d_{m-1}$, and so by induction $K\mathscr{G}|_{X}$ is regular. If we can show that $K\mathscr{G}|_{U}$ is regular, then Proposition 2.1 will imply that $K\mathscr{G}$ is regular.
Let $V\subseteq U$ be a compact open set. Then $V$ is clopen in $\mathscr{G}^{(0)}$ and hence $\mathop{\boldsymbol{r}}^{-1}(V)\subseteq\mathscr{G}$ is clopen, hence quasi-compact and open. Thus $W=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}^{-1}(V))$ is a compact open invariant subset of $U$ containing $V$. Hence the compact open invariant subsets of $U$ are cofinal among compact open subsets of $U$. Moreover, if $W$ is a compact open invariant subspace of $U$, then $W$ is clopen in $\mathscr{G}^{(0)}$ (as the latter is Hausdorff), and so $\mathscr{G}|_{W}=\mathop{\boldsymbol{r}}^{-1}(W)$ is a clopen subgroupoid of $\mathscr{G}$, hence quasi-compact, with all orbits of size $d_{m}$ (as $W\subseteq U$ is invariant). Hence $K\mathscr{G}|_{W}$ is regular by Lemma 3.10. Since $K\mathscr{G}|_{U}$ is the direct limit of the $K\mathscr{G}|_{W}$ where $W$ runs over the compact open invariant subsets of $U$, we conclude that $K\mathscr{G}|_{U}$ is regular from Proposition 2.1. This completes the proof. ∎
@@
If one follows the proofs carefully, it is not difficult to see that if $K$ is a field of characteristic $p>0$ and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that no isotropy element has order divisible by $p$, then $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple algebras, and hence if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial. This just uses that the class under consideration is closed under matrix amplification and if an algebra has a finite ideal filtration with factors of this form, then the algebra has this form.

### v3

@@ 2.4. Regular rings
Let $U\subseteq\mathscr{G}^{(0)}$ be the identity of a maximal subgroup $H$, and note that $H$ consists of those $V\in S$ with $\mathop{\boldsymbol{s}}(V)=U=\mathop{\boldsymbol{r}}\nolimits(V)$. If $V\in H$ and $x\in U$, then denote by $v_{x}$ the unique element of $V$ with $\mathop{\boldsymbol{r}}\nolimits(v_{x})=x$. Since $\mathscr{G}$ consists of isotropy, $\mathop{\boldsymbol{s}}(v_{x})=x$ as well. Then we have an injective group homomorphism $\psi\colon H\to\prod_{x\in U}\mathscr{G}_{x}^{x}$ given by $\psi(V)_{x}=v_{x}$. Note that if $V\in H$ has order $p$, then $\psi(V)$ has order $p$, and hence $v_{x}^{p}=x$ for all $x\in U$, and some $v_{x}$ has order $p$. Therefore, $p$ is invertible in $R$ by assumption. ∎
@@
Let $\mathscr{G}$ be a quasi-compact ample groupoid and $R$ a regular unital ring such that the order of any cyclic subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$. Suppose that all orbits of $\mathscr{G}$ have the same cardinality. Then $R\mathscr{G}$ is regular.
@@
It follows from Lemma 3.3 that the size of each orbit is finite, say that this common size is $n$. Choose, for each $x\in\mathscr{G}^{(0)}$, elements $\gamma_{x,1},\ldots,\gamma_{x,n}$ with $\mathop{\boldsymbol{s}}(\gamma_{x,i})=x$ and $\mathcal{O}_{x}=\{\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,n})\}$. We may assume without loss of generality that $\gamma_{x,1}=x$ (the identity at $x$). Choose for each $i$ a compact open bisection $U^{x}_{i,1}$ containing $\gamma_{x,i}$. We may assume that $U^{x}_{1,1}\subseteq\mathscr{G}^{(0)}$. Since $\mathscr{G}^{(0)}$ is Hausdorff with a basis of compact open sets and $\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(\gamma_{x_{n}})$ are pairwise distinct, we can find pairwise disjoint compact open neighborhoods $V_{1},\ldots,V_{n}$ of $\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,n})$, respectively, in $\mathscr{G}^{(0)}$. Replacing $U^{x}_{i,1}$ by $V_{i}U^{x}_{i,1}$, we may assume that the sets $U^{x}_{i}=\mathop{\boldsymbol{r}}\nolimits(U^{x}_{i,1})$ are pairwise disjoint. Also, putting $U=\bigcap_{i=1}^{n}\mathop{\boldsymbol{s}}(U^{x}_{i,1})$ and replacing $U^{x}_{i,1}$ by $U^{x}_{i,1}U$, we may assume that $\mathop{\boldsymbol{s}}(U^{x}_{i,1})=\mathop{\boldsymbol{s}}(U^{x}_{1,1})=U^{x}_{1}$ for all $i$. Put $U^{x}_{i,j}=U^{x}_{i,1}(U^{x}_{j,1})^{-1}$. Note that $\mathop{\boldsymbol{s}}(U^{x}_{i,j})=U^{x}_{j}$ and $\mathop{\boldsymbol{r}}\nolimits(U^{x}_{i,j})=U^{x}_{i}$ for all $i,j$. Also note that $U^{x}_{i,i}=U^{x}_{i}$ for $i=1,\ldots,n$. Let $U^{x}=\bigcup_{i=1}^{n}U^{x}_{i}$. Then $U^{x}$ is compact open and we claim that it is invariant. For suppose that $y\in U^{x}_{j}$. Then since $U^{x}_{j}=\mathop{\boldsymbol{s}}(U^{x}_{i,j})$ and $U^{x}_{i}=\mathop{\boldsymbol{r}}\nolimits(U^{x}_{i,j})$ for all $i=1,\ldots,n$, we can find elements of the orbit of $y$ in each of $U^{x}_{1},\ldots,U^{x}_{n}$. Since these sets are pairwise disjoint, this yields $n$ elements in the orbit of $y$ and hence the whole orbit by assumption. Thus each $U^{x}$ is invariant and each $U^{x}_{i}$ contains exactly one element from each orbit of $U^{x}$. Also note that
@@
Another observation is that the isotropy subgroupoid $\mathop{\mathrm{Iso}}(\mathscr{G})$ of $\mathscr{G}$ is clopen, and hence quasi-compact. It is closed since $\mathscr{G}^{(0)}$ is Hausdorff and it is the equalizer of $\mathop{\boldsymbol{s}},\mathop{\boldsymbol{r}}\nolimits$. It is open because if $\gamma\colon x\to x$ is isotropy and $V$ is a compact open bisection containing $\gamma$, then $U^{x}_{1}VU^{x}_{1}$ is a compact open bisection containing $\gamma$ and consisting of isotropy as $x\in U^{x}_{1}$, and $U^{x}_{1}$ contains exactly only one element from each orbit of the invariant compact open set $U^{x}$, whence $\mathop{\boldsymbol{s}}(\alpha),\mathop{\boldsymbol{r}}\nolimits(\alpha)\in U^{x}_{1}$ implies $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}\nolimits(\alpha)$.
Since $\mathscr{G}^{(0)}$ is compact and the $U^{x}$ cover $\mathscr{G}^{(0)}$, we can find finitely may points $x_{1},\ldots,x_{r}$ such that $\mathscr{G}^{(0)}=U^{x_{1}}\cup\cdots\cup U^{x_{r}}$. Without loss of generality, we may assume that $U^{x_{i}}\nsubseteq\bigcup_{j\neq i}U^{x_{j}}$ for all $i$. Put $V_{1}=U^{x_{1}},V_{2}=U^{x_{2}}\setminus V_{1},\ldots,V_{r}=U^{x_{r}}\setminus\bigcup_{i=1}^{r-1}V_{i}$. Then $V_{1},\ldots,V_{r}$ is partition of $\mathscr{G}^{(0)}$ into compact open invariant sets. In particular note that $V_{i}U=UV_{i}$ for any compact open bisection $U$. It follows that if $\mathscr{G}_{i}=\mathscr{G}|_{V_{i}}$, then $R\mathscr{G}\cong\prod_{k=1}^{r}R\mathscr{G}_{k}$ and $R\mathscr{G}_{k}=1_{V_{k}}R\mathscr{G}1_{V_{k}}$.
Note that since $\mathscr{G}_{k}=\mathop{\boldsymbol{r}}\nolimits^{-1}(V_{k})$ is clopen in $\mathscr{G}$ it is quasi-compact. Also $\mathop{\mathrm{Iso}}(\mathscr{G}_{k})=\mathop{\mathrm{Iso}}(\mathscr{G})\cap\mathscr{G}_{k}$ is clopen and hence quasi-compact. Put $U^{k}_{i,j}=V_{k}U^{x_{k}}_{i,j}=U^{x_{k}}_{i,j}V_{k}$; this is a compact open bisection in $\mathscr{G}_{k}$. Also note that $U^{k}_{1,1}\subseteq U^{x_{k}}_{1}$ does not contain any two points in the same orbit by the previous discussion. Hence $U^{k}_{1,1}\mathscr{G}_{k}U^{k}_{1,1}=\mathscr{G}_{k}|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})\cap\mathop{\boldsymbol{r}}\nolimits^{-1}(U^{k}_{1,1})$ is clopen (as $U^{k}_{1,1}$ is a clopen subspace of $\mathscr{G}^{(0)}$) and thus a quasi-compact group bundle. Note that $V_{k}=V_{k}U^{x_{k}}=\bigcup_{i=1}^{n}U^{k}_{i,i}$ and hence it follows from (3.2) that the $1_{U^{k}_{i,j}}$ form a set of matrix units for $R\mathscr{G}_{k}$. Thus $R\mathscr{G}_{k}\cong M_{n}(1_{U^{k}_{1,1}}R\mathscr{G}_{k}1_{U^{k}_{1,1}})=M_{n}(R\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}})$. We deduce that $R\mathscr{G}$ is regular from Propositions 2.1 and 3.7. ∎
@@
Clearly $O_{k}$ is invariant; we claim that it is open. Suppose that $x\in O_{k}$. Let $x=y_{1},y_{2}\ldots,y_{k}$ be $k$ distinct elements in the orbit of $x$. As in the proof of Lemma 3.8, we can find open bisections $U_{1},\ldots,U_{k}$ with $x\in\mathop{\boldsymbol{s}}(U_{i})$, $y_{i}\in\mathop{\boldsymbol{r}}\nolimits(U_{i})$, for $i=1,\ldots,k$, and $\mathop{\boldsymbol{r}}\nolimits(U_{1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(U_{k})$ pairwise disjoint (using that $\mathscr{G}^{(0)}$ is Hausdorff). Then $U=\bigcap_{i=1}^{k}\mathop{\boldsymbol{s}}(U_{i})$ is a neighborhood of $x$ and if $y\in U$, then each of $\mathop{\boldsymbol{r}}\nolimits(U_{1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(U_{k})$ contains an element of the orbit of $y$. Since these sets are pairwise disjoint, we conclude that $y\in O_{k}$. ∎
We can now prove the sufficiency of the conditions in Theorem A.
Proposition 3.10.
Suppose that $R$ is a regular unital ring and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that the order any finite subgroup of an isotropy group is invertible in $R$. Then $R\mathscr{G}$ is regular.
@@
Let $d_{1}<d_{2}<\cdots<d_{m}$ be the distinct orbit sizes. We prove the result by induction on $m$. If $m=1$, then all orbits of $\mathscr{G}$ have the same size and the result follows from Lemma 3.8. Let $U=\{x\in\mathscr{G}^{(0)}\mid|\mathcal{O}_{x}|=d_{m}\}$. Then $U$ is open and invariant by Lemma 3.9 once we note that $U=O_{d_{m}}$ in the notation of that lemma. Let $X=\mathscr{G}^{(0)}\setminus U$. Then $I=R\mathscr{G}|_{U}$ is an ideal of $R\mathscr{G}$ and $R\mathscr{G}/I\cong R\mathscr{G}|_{X}$. Note that $\mathscr{G}|_{U}=\mathop{\boldsymbol{r}}\nolimits^{-1}(U)$ is open and $\mathscr{G}|_{X}=\mathop{\boldsymbol{r}}\nolimits^{-1}(X)$ is closed and hence quasi-compact. By construction, the orbits sizes in $\mathscr{G}|_{X}$ are $d_{1}<d_{2}<\cdots<d_{m-1}$, and so by induction $R\mathscr{G}|_{X}$ is regular. If we can show that $R\mathscr{G}|_{U}$ is regular, then Proposition 2.1 will imply that $R\mathscr{G}$ is regular.
Let $V\subseteq U$ be a compact open set. Then $V$ is clopen in $\mathscr{G}^{(0)}$, and hence $\mathop{\boldsymbol{r}}\nolimits^{-1}(V)\subseteq\mathscr{G}$ is clopen, whence quasi-compact and open. Thus $W=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}\nolimits^{-1}(V))$ is a compact open invariant subset of $U$ containing $V$. Hence the compact open invariant subsets of $U$ are cofinal among compact open subsets of $U$. Moreover, if $W$ is a compact open invariant subspace of $U$, then $W$ is clopen in $\mathscr{G}^{(0)}$ (as the latter is Hausdorff), and so $\mathscr{G}|_{W}=\mathop{\boldsymbol{r}}\nolimits^{-1}(W)$ is a clopen subgroupoid of $\mathscr{G}$, hence quasi-compact, with all orbits of size $d_{m}$ (as $W\subseteq U$ is invariant). Hence $R\mathscr{G}|_{W}$ is regular by Lemma 3.8. Since $R\mathscr{G}|_{U}$ is the direct limit of the $R\mathscr{G}|_{W}$ where $W$ runs over the compact open invariant subsets of $U$, we conclude that $R\mathscr{G}|_{U}$ is regular from Proposition 2.1. This completes the proof. ∎
@@
If one follows the proofs carefully, it is not difficult to see that if $K$ is a field and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that no isotropy element has order divisible by the characteristic of $K$, then $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple algebras, and hence if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial. This just uses that the class under consideration is closed under matrix amplification and if an algebra has a finite ideal filtration with factors of this form, then the algebra has this form.
We may now prove Theorem A.
Proof of Theorem A.
Necessity of (1)–(3) follows from Propositions 3.1 and 3.6. Sufficiency of these conditions follows from Proposition 3.10. ∎